OpenStudy (kaloemalia):

i think i have the answer but i'm not sure.. What is the slope of the line? Y=2x-7 I think the slope is -2.. is that correct?

OpenStudy (dtan5457):

Hmm, why -2?

OpenStudy (dtan5457):

(wrong sign)

OpenStudy (dtan5457):

y=mx+b Y=2x-7 m=slope

OpenStudy (anonymous):

Y=mx + b M is the slope B is the y intercept

OpenStudy (kaloemalia):

so it is just 2?

OpenStudy (dtan5457):

Yeah

OpenStudy (anonymous):

Y=mx+b M is the slope

OpenStudy (anonymous):

right.

OpenStudy (kaloemalia):

okay thanks!